### v1

@@ E.2 Evaluation Config, Manifest & Report
Metrics are declared under metrics with their own clients and prompts shown in E.1, allowing the judge to differ from the proxy/assistant. We show two metrics, GTEval and Pairwise Indistinguishability, both routed to Claude-4-Sonnet at temperature=0. Each metric includes num_judge_samples (for repeated scoring) and compute_controls=true, which triggers the HH/PP control comparisons used for calibration and bias checks. Finally, the task_drivers section binds the dataset to the Mirror Conversation Driver, which synthesizes multi-turn dialogues by alternating the user-proxy and assistant models; here, the assistant is also GPT-4o at temperature=0, ensuring low-variance responses for per-proxy comparability. Together, these blocks yield a self-contained manifest that our planner expands into (proxy, dataset, metric, seed) units, enabling exact replay and variance-aware aggregation across runs. We illustrate an example of such manifest file in Figure 16, which is a manifest produced using config shown in Figure 15.

### v2

@@ B.2. Evaluation Config, Manifest & Report
Metrics are declared under metrics with their own clients and prompts shown in B.1, allowing the judge to differ from the proxy/assistant. We show two metrics, GTEval and Pairwise Indistinguishability, both routed to Claude-4-Sonnet at temperature=0. Each metric includes num_judge_samples (for repeated scoring) and compute_controls=true, which triggers the HH/PP control comparisons used for calibration and bias checks. Finally, the task drivers section binds the dataset to the Mirror Conversation Driver, which synthesizes multi-turn dialogues by alternating the user-proxy and assistant models; here, the assistant is also GPT-4o at temperature=0, ensuring low-variance responses for per-proxy comparability. Together, these blocks yield a self-contained manifest that our planner expands into (proxy, dataset, metric, seed) units, enabling exact replay and variance-aware aggregation across runs. We illustrate an example of such manifest file in Figure 17, which is a manifest produced using config shown in Figure 19.